OpenStudy (anonymous):

Help A ball is thrown upward from the top of a building. The function below shows the height of the ball above the front, f(t), in feet, at different times (t) in seconds: f(t) = = -16t^2 + 34t + 80

OpenStudy (anonymous):

the average rate of change of f(x) from x = 5 to x= 7is_____.

OpenStudy (ddcamp):

\[Average\ Rate\ of\ Change = \frac{Change\ of\ x}{Change\ of\ time}\]

OpenStudy (ddcamp):

Or \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

I was thinking i needed to plug in the numbers into the orignal equation?

OpenStudy (ddcamp):

Yeah. Plug 7 and 5 into the equation to find f(7) and f(5). Subtract f(5) from f(7), then divide by 2 (=7-5)

OpenStudy (anonymous):

Dang there are some crazy numbers im getting? should i be getting like 400 and 170 for the numbers in x = 5

OpenStudy (anonymous):

f(5) = -150 now should i divide by 2?